# Investigating Popularity Bias Amplification in Recommender Systems Employed in the Entertainment Domain

### v1

@@ Abstract.
Recommender systems have become an integral part of our daily online experience by analyzing past user behavior to suggest relevant content in entertainment domains such as music, movies, and books. Today, they are among the most widely used applications of AI and machine learning. Consequently, regulations and guidelines for trustworthy AI, such as the European AI Act, which addresses issues like bias and fairness, are highly relevant to the design, development, and evaluation of recommender systems. One particularly important type of bias in this context is popularity bias, which results in the unfair underrepresentation of less popular content in recommendation lists. This work summarizes our research on investigating the amplification of popularity bias in recommender systems within the entertainment sector. Analyzing datasets from three entertainment domains, music, movies, and anime, we demonstrate that an item’s recommendation frequency is positively correlated with its popularity. As a result, user groups with little interest in popular content receive less accurate recommendations compared to those who prefer widely popular items. Furthermore, we aim to better understand the connection between recommendation accuracy, calibration quality of algorithms, and popularity bias amplification.
@@ 1. Introduction and Motivation
Recommender systems have become one of the most prevalent applications of machine learning and AI, shaping our daily online experiences. They play a crucial role in helping users navigate vast and complex information spaces by identifying relevant content (Burke et al., 2011; Jannach et al., 2016). Since their early implementations (Resnick et al., 1994), these systems have relied on analyzing past user behavior to construct user models and provide recommendations—not only for items like movies, music, and books but also for social connections in online networks (Kowald et al., 2013; Lacic et al., 2015; Eirinaki et al., 2018).
@@
The present work summarizes our research on investigating the amplification of popularity bias in recommender systems within the entertainment domain. Analyzing datasets from three sectors, music, movies, and animes, we demonstrate that popularity bias disproportionately impacts user groups with little interest in popular content, as they receive less accurate recommendations compared to those who favor popular items (Kowald et al., 2020, 2021; Kowald and Lacic, 2022; Kowald et al., 2023). Furthermore, we seek to gain deeper insights into the relationship between recommendation accuracy, algorithmic calibration quality, and the amplification of popularity bias. Finally, we propose directions for future research.
@@ 4. Conclusion and Future Research Directions
In this work, we summarize our research on investigating the amplification of popularity bias in recommender systems within the entertainment sector. Analyzing datasets from music, movies, and anime, we show that recommendation frequency increases with item popularity, disadvantaging users with little interest in popular content. Additionally, we explore the interplay between recommendation accuracy, algorithmic calibration, and the amplification of popularity bias. Future research should investigate additional domains with respect to popularity bias amplification, and propose robust methods for mitigating popularity bias to foster fair recommendations.

### v2

@@ Abstract.
Recommender systems have become an integral part of our daily online experience by analyzing past user behavior to suggest relevant content in entertainment domains such as music, movies, and books. Today, they are among the most widely used applications of AI and machine learning. Consequently, regulations and guidelines for trustworthy AI, such as the European AI Act, which addresses issues like bias and fairness, are highly relevant to the design, development, and evaluation of recommender systems. One particularly important type of bias in this context is popularity bias, which results in the unfair underrepresentation of less popular content in recommendation lists. This work summarizes our research on investigating the amplification of popularity bias in recommender systems within the entertainment sector. Analyzing datasets from three entertainment domains, music, movies, and anime, we demonstrate that an item’s recommendation frequency is positively correlated with its popularity. As a result, user groups with little interest in popular content receive less accurate recommendations compared to those who prefer widely popular items. Furthermore, this work contributes to a better understanding of the connection between recommendation accuracy, calibration quality of algorithms, and popularity bias amplification.
@@ 1. Introduction and Motivation
Recommender systems have become one of the most prevalent applications of machine learning and AI, shaping our daily online experiences. They play a crucial role in helping users navigate vast and complex information spaces by identifying relevant content (Burke et al., 2011; Jannach et al., 2016). Since their early implementations (Resnick et al., 1994), these systems have relied on analyzing past user behavior to construct user models and provide recommendations, not only for items like movies, music, and books, but also for social connections in online networks (Kowald et al., 2013; Lacic et al., 2015; Eirinaki et al., 2018).
@@
The present work summarizes our research on investigating the amplification of popularity bias in recommender systems within the entertainment domain. Analyzing datasets from three sectors, music, movies, and animes, we demonstrate that popularity bias disproportionately impacts user groups with little interest in popular content, as they receive less accurate recommendations compared to those who favor popular items (Kowald et al., 2020, 2021; Kowald and Lacic, 2022; Kowald et al., 2023). Furthermore, this work provides deeper insights into the relationship between recommendation accuracy, algorithmic calibration quality, and the amplification of popularity bias. Finally, we propose directions for future research.
@@ 4. Conclusion and Future Research Directions
In this work, we summarize our research on investigating the amplification of popularity bias in recommender systems within the entertainment sector. Analyzing datasets from music, movies, and anime, we show that recommendation frequency increases with item popularity, disadvantaging users with little interest in popular content. Additionally, we explore the interplay between recommendation accuracy, algorithmic calibration, and the amplification of popularity bias. Future research should investigate additional domains with respect to popularity bias amplification (e.g., news (Lacic et al., 2022)), and propose robust methods for mitigating popularity bias (e.g., (Klimashevskaia et al., 2022)).
Acknowledgements. This work was supported by the FFG COMET program. A full version of this article was accepted as a habilitation at Graz University of Technology in June 2024, and is available via Arxiv (Kowald, 2024).